When resistance of 2 ohm is connected across the terminals of the cell the current is 0.5 A and when the resistance is increased to 5 ohm the current is 0.25 A. The emf of the cell is ?


1
Expert's answer
2021-06-21T16:25:37-0400

Gives

Resistance

"R=2\Omega"

Current

"I=0.5A"

Electromotive force

Emf="I(r+R)\rightarrow(1)"

Emf of the cell=V

"V=I(r+R)"

Pu value

V=0.5(2+r)"\rightarrow(2)"

When resistance increase

R=5"\Omega"

"I=0.25A"

V=0.25(5+r)"\rightarrow(3)"

equation (2)and(3)are equal

Above both term potential is equal

0.25(5+r)=0.5(2+r)

(5+r)=2(2+r)

r=1"\Omega"

We know that

V=0.25(5+r)

Put r value

V=0.25(5+1)=1.5V

e.m.fof thr cell

V=1.5V